1. Automotive air bags inflate when a sample of sodium azide, NaN3, is very rapidly decomposed: 2NaN3 (s) ―—› 2Na(s) + 3N2(g) What mass of sodium azide is required to produce 2.6 ft3 (73.6 L) of nitrogen gas with a density of 1.25 g/L?
Answer #1

mass of N2 produced = density * volume
= 1.25 g/L * 73.6 L
= 92 g

number of moles of N2 produced = mass/molar mass
= 92/28
= 3.286 mol

moles of NaN3 required = (2/3)*moles of N2
= (2/3)*3.286 mol
=2.19 mol

mass of NaN3 required= number of mol *molar mass of NaN3
= 2.19*65
= 142.4 g
Answer: 142.4 g
Answer #2